# Inflation with the standard and Randall–Sundrum model in the Two-time Physics

### 16 Mar 2026 v1

@@ Abstract
We propose a scalar inflationary potential as $V(\phi)=M^{4}\phi^{2n-2}(\phi^{2n}+m^{2n})^{1/n-1}$. This potetial similar to the shaft inflation one. The potential may come from the Higgs-dilaton potential in the Two-time (2T) physics, especially in the case where $n=3$, this suggests an explanation for the inflationary potential. Therefore, we call it shaft-warm inflation potential for short. The slow-roll scenario is recomputed in the 4-dimension (4D) and Randall-Sundrum II (RSII) frameworks. The tensor-to-scalar ratio in RSII is always higher than in 4D and is in good agreement with the experimental data of BICEP2 and Planck. When compared with Planck data we estimate $M_{5}$ to be around $1-2]\times 10^{16}$ GeV. Furthermore, the potential allows much lower scalar field exponents than other potentials, which results in high agreement with experimental data.
@@ II.1 The 2T metric reduction
We make a final change of variable
@@ III The Shaft-warm inflation in the 4D spacetime
When $\phi\ll m$ the potential becomes monomial, but when $n\gg m$ the potential with a constant form or superpotential will be linear in $\phi$. These are similar to those in the shaft inflation.
@@
With $N\sim 50-60$, the running spectral index is about $\text{a few}\times 10^{-14}$, is very small, thus it is compatible with observations of the Planck Collaboration [18]. These results are better than other models. However, $r$ in Eq.28 is too small compared to observations. This is also the reason why we consider the inflation scenario in RSII model. This will be explained in the section V.
@@ IV The Shaft-warm inflation in the RSII model
The tensor-to-scalar ratio is proportional to $\epsilon$, we will derive it as a function of $N$ using Eq. 38,
@@ VI Conclusions and discussions
The 2T model contains the dilaton field, which gives us a natural explanation of the dilaton field in the context of gravity, and the dilaton as a candidate for the Inflaton.

### 07 May 2026 v2

@@ Abstract
We propose a scalar inflationary potential as $V(\phi)=M^{4}\phi^{2n-2}(\phi^{2n}+m^{2n})^{1/n-1}$. This potetial similar to the shaft inflation one. The potential may come from the Higgs-dilaton potential in the Two-time (2T) physics, especially in the case where $n=3$, this suggests an explanation for the inflationary potential. Therefore, we call it the shaft-warm inflation potential for short. The slow-roll scenario is recomputed in the 4-dimension (4D) and Randall-Sundrum II (RSII) frameworks. The tensor-to-scalar ratio in RSII is always higher than in 4D and is in good agreement with the experimental data of BICEP2 and Planck. When compared with Planck data we estimate $M_{5}$ to be around $1-2]\times 10^{16}$ GeV. Furthermore, the potential allows much lower scalar field exponents than other potentials, which results in high agreement with experimental data.
@@ II.1 The 2T metric reduction
We make a final change of variable,
@@ III The Shaft-warm inflation in the 4D spacetime
When $\phi\ll m$ the potential becomes monomial, but when $\phi\gg m$ the potential with a constant form or superpotential will be linear in $\phi$. These are similar to those in the shaft inflation.
@@
With $N\sim 50-60$, the running spectral index is about $\text{a few}\times 10^{-4}$, is very small, thus it is compatible with observations of the Planck Collaboration [18]. These results are better than other models. However, $r$ in Eq.28 is too small compared to observations. This is also the reason why we consider the inflation scenario in RSII model. This will be explained in the section V.
@@ IV The Shaft-warm inflation in the RSII model
By using Eq. 38, the tensor-to-scalar ratio is proportional to $\epsilon$, we will derive it as a function of $N$,
@@ VI Conclusions and discussions
The 2T model contains the dilaton field, which gives us a natural explanation of the dilaton field in the context of gravity, and the dilaton can be a candidate for the inflaton.